OpenStudy (wessicawynnn):

Find the quotient. h 6 ÷ h

OpenStudy (phi):

if this is \[ \frac{6h}{h} \] then that is the same as \[ 6 \cdot \frac{h}{h} \] now remember anything divided by itself is 1 (that works for letters as well as numbers)